OpenStudy (anonymous):

I GIVE MEDALS AND FAN!!!!!!!!!!!!!!!!!! Solve for t. 10,000= 39,145 (1+0.03/ 12)^12T

OpenStudy (johnweldon1993):

Lets start by dividing both sides by 39145

OpenStudy (anonymous):

and then one side cancels out right?

OpenStudy (johnweldon1993):

Well the 39145 on the right cancels out yes...so what do you have after you do that?

OpenStudy (anonymous):

divide 39145 by 10000

OpenStudy (johnweldon1993):

Other way around...divide 10000 by 39145

OpenStudy (anonymous):

i got 0.255460467. would i round ?

OpenStudy (johnweldon1993):

\[\large \frac{10000}{39145} =\cancel{\frac{39145}{39145}}(1 + \frac{0.03}{12})^{12T}\] Yeah we can round that to .255 for now :) so we now have \[\large .255 = (1 + \frac{0.03}{12})^{12T}\]

OpenStudy (anonymous):

((((240.743759253/6)*2)+20)-100)=your answer, (simplified it/hid it, you still have to work for it)

OpenStudy (anonymous):

don't use google, it rounds too much

OpenStudy (anonymous):

do i divide 0.03 and 12 ?

OpenStudy (johnweldon1993):

Now lets simplfy that stuff in the parenthesis... What is .003 / 12 ? and then add 1

OpenStudy (johnweldon1993):

Sorry 0.03 / 12

OpenStudy (anonymous):

i got 1.0025

OpenStudy (anonymous):

its less than 0.5 t<0.5

OpenStudy (johnweldon1993):

Perfect...so now we have \[\large .255 = (1.0025)^{12T}\] Now...we insert Logarithms...have you worked with them before?

OpenStudy (anonymous):

yes but i dont understand them...

OpenStudy (johnweldon1993):

That okay...so 1 rule to remember *that we will use here* is \[\large lna^x = x \times lna\] meaning when you take the og of something with a variable as the exponent......that exponent comes down and multiplies to the log So here... \[\large ln(.255) = 12T \times ln(1.0025)\]

OpenStudy (anonymous):

so i would find the log of .255 & 1.0025?

OpenStudy (johnweldon1993):

Not quite...now you divide both sides by ln(1.0025) because we want to isolate that 12T right? so \[\large \frac{ln(.255)}{ln(1.0025)} = 12T\] Right?

OpenStudy (anonymous):

so now i can find the log or do i didvide them?

OpenStudy (johnweldon1993):

Well both ...unless you have a calculator of course then just input them :)

OpenStudy (anonymous):

okay i got -547.2796551 o:

OpenStudy (johnweldon1993):

Haha don't freak out...that's correct :) Just 1 more step though... \[\large -547.2796551 = 12T\] To solve for 't' we divide both sides by 12....so T = ?

OpenStudy (anonymous):

LOL okay c: T=-45.60663793

OpenStudy (johnweldon1993):

And there is your answer :) * I checked it it's right ;P

OpenStudy (anonymous):

Thanks!

OpenStudy (johnweldon1993):

Hope that all made sense hun :)